OpenStudy (stuck-help):

- The lengths of the sides of a sandbox are 9ft,6ft, and N ft. write a compound inequality that describes all possible lengths of N

OpenStudy (stuck-help):

@Missiey

OpenStudy (qwertty123):

hello, Hold on while I try to explain this.

OpenStudy (triciaal):

so the sandbox is triangular

OpenStudy (qwertty123):

computer is slow....

OpenStudy (stuck-help):

@Missiey can you help

OpenStudy (qwertty123):

it sounds like the "box" has only 3 sides... a triangle. the third side of a triangle has to be shorter than the sum of the other two sides and bigger than the difference of the other two sides. can you figure out the longest n can be? and the shortest n can be ?

OpenStudy (stuck-help):

the answer would be n = less than 15 and more than 3

OpenStudy (qwertty123):

nope Side 1 is 9 and side 2 is 6 so 9-6 =? and 9+6=?

OpenStudy (stuck-help):

9-6=3 and 9+6= 15 so less than 15 and more than 3

OpenStudy (qwertty123):

Yes!

OpenStudy (stuck-help):

so it was right because that is exactly what i said at first

OpenStudy (qwertty123):

that means the third side (n) has to be between those two numbers. Can you write that down ? example 3 < n is part of the answer

OpenStudy (qwertty123):

@stuck-help

OpenStudy (qwertty123):

@stuck-help what is your answer?